in a chemistry class 5 liters of a 4% saline solution must be mixed with a 10% solution to get a 6% solution how many liters of the 10% solution are needed
Answer by stanbon(75887)

in a chemistry class 5 liters of a 4% saline solution must be mixed with a 10% solution to get a 6% solution how many liters of the 10% solution are needed
-------------------------
Equation:
salt + salt = salt
0.04*5 + 0.10x = 0.06(5+x)
---
Multiply thru by 100 to get:
4*5 + 10x = 6*5+6x
4x = 2*5
x = 2.5 liters (amt. of 10% solution needed for the mixture)
========================================
Cheers,
Stan H.